A light hits a surface at an angle of 16°. The material on the other side of the surface has an index of refraction of 1.44 and the light is making an angle of 15° with the surface when it exits on this side. What is the index of refraction on the incident side of the interface?

1
Expert's answer
2022-05-14T16:45:58-0400

"n_2=n_1\frac{\sin\alpha}{\sin\beta}=1.53."